OpenStudy (anonymous):

what is the positive numbers for the solution set of |-2x-4|<=12

OpenStudy (anonymous):

-12 <= -2x-4 <= 12

OpenStudy (anonymous):

if i solve this then i get 4<=x<= -8

OpenStudy (anonymous):

everyone must have the same homework. this is the 5th time i have seen this question

OpenStudy (anonymous):

@gymgirl no your answer is wrong

OpenStudy (anonymous):

there is no such thing as \[4<x\leq-8\]

OpenStudy (anonymous):

first to make the problem easier note that \[|-2x-4|=|2x+4| \]and then solve \[|2x+4|<12\] via the compound inequality \[-12<2x+4<12\] \[-16<2x<8\] \[-8<x<4\]

OpenStudy (anonymous):

You forgot the equal sign

OpenStudy (anonymous):

if you want to skip the first step we can redo the problem as \[|-2x-4|<12\] \[-12<-2x-4<12\] \[-8<-2x<16\] and now when we divide by -2 we have to change both inequalities because -2 is negative to get \[4>x>-8\] or \[-8<x<4\]

OpenStudy (anonymous):

the problem should be less than or equal to

OpenStudy (anonymous):

-8<=x<=4

OpenStudy (anonymous):

@heromiles you are right of course. it should be \[-8\leq x\leq4\] but method is identical

OpenStudy (anonymous):

ok i got that as my answer. but i need to know that solution set and it has to be positive numbers. that is what i am confused about